OpenStudy (anonymous):

x + 5y = -45 -2x - 15y = 140 elimination method

OpenStudy (anonymous):

x + 5y = -45 5y=-45-x y=(-45-x)/5

OpenStudy (anonymous):

y=-10 x=5

OpenStudy (anonymous):

substitute this y in second equation u'll get the answer

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

x=5 y=-10

OpenStudy (anonymous):

@gillam1021 that's the answer

OpenStudy (anonymous):

@gillam1021 please don't ask for answers...this is open study not answers.com please read the code of conduct if you haven't http://openstudy.com/code-of-conduct